# Test-time Correction with Human Feedback: An Online 3D Detection System via Visual Prompting

### v2

@@ 3.3 Visual Prompt Buffer
Visual prompt buffer is a queue that stores user-provided visual prompts of missed objects during online inference. In the default setting, these visual prompts are the 2D output regions from the model in previous frames, based on the user’s point or box prompts at that time. The flexibility of the TTC system also allows users to freely select visual prompts, either image contents from the current scene or customized objects from the Internet. This versatility makes the TTC system applicable to a wide spectrum of scenarios.
Dequeue. To prevent the buffer from growing indefinitely, we design a “dequeue” mechanism. We filter out visual prompts with low confidence, which likely no longer appear in the scene. We also use the intersection-over-union (IoU) between predictions to identify and remove redundant visual prompts. This dynamic update and maintenance ensure a balance of TTC system between latency and accuracy by adaptively incorporating and pruning feedback online.
@@ 4 Experiments
We now proceed to evaluate our TTC system from two key aspects:
@@ 4.1 Experimental Setup
Dataset. Experiments are done on nuScenes dataset [1] with 1,000 autonomous driving sequences, one of the most popular datasets for autonomous driving research.
@@
Tasks. Experiments are conducted in two aspects to answer the questions above. The first experiment aims to test the TTC system in boosting the performance of offline-trained 3D detectors via online error rectification. We conduct this verification by applying TTC system to various established offline 3D detectors [67, 27, 62, 32, 54, 31, 57].
@@
Then, we validate the TTC system to correct detection errors in out-of-training-distribution scenarios. To conduct a thorough quantitative analysis of this aspect, we established four tasks under different settings: (a.) discarding 80% labels of distant objects farther than 30m during training, and test the error corrections for detecting distant objects; (b.) discarding 80% labels of instances labeled as vehicle, including “car”, “truck”, “C.V.”, “bus”, and “trailer”, and test the error correction in vehicles; (c.) discarding all annotations of class “truck” and “bus”, and test the zero-shot ability of TTC on those discarded classes; (d.) discarding all training data of the scenario of “Nighttime” and “Rainy”, and test the improvements of TTC on scenarios with domain gap. For these tasks, we base TTC system on the monocular algorithm, MonoDETR [67], and test it under each setting separately, as the monocular setting demonstrates the most general applicability. Based on TTC-MonoDETR, we also present qualitative results to show the potential of TTC to address corner cases in challenging real-world scenarios through human feedback during test time. Through these experiments, we demonstrate TTC as an effective system to adapt offline 3D detectors to challenging scenarios.
Entity Detection Score.
@@
As we focus on enabling existing 3D detectors to rectify missing detections, both in-distribution and out-of-distribution, we adopt a class-agnostic setting. We remove all class annotations of 3D objects during training, treating all objects as entities [24, 45]. For evaluation, we use the Entity Detection Score (EDS), a class-agnostic version of the nuScenes Detection Score (NDS) that emphasizes recall. A detailed analysis of the rationale behind EDS can be found in Appendix Section C.3.
Implementation Details. We implement our method based on MMDetection3D codebase [8], and conduct all experiments on a server with 8$\times$ A100 GPUs. In OA module, we use a ResNet18 [17] to extract visual prompt features. All visual prompts are resized to $224\times 224$ before being sent to OA module. For training, we use AdamW [23, 36] with a batch size of 16. We initialize the learning rate as 2e-4, adjusted by the cosine annealing policy. When training point and box prompts, we simulate user inputs with noise by adding perturbations to the ground truth. To ensure the visual prompts are robust across multiple scenes, timestamps, and styles, we choose visual prompts of target objects not only from the image patches of the current frame but also randomly from previous and future frames within a range of ±$\pm 5$. Flip operations are used as data augmentations. During testing, we mimic user intervention. We simulate the user’s online missing feedback by comparing the distance between ground truth and detected 3D boxes. Ground-truths without any detection within 2m are considered missed and added into the prompt buffer. We remove redundant predictions by non-max-suppression (NMS) with the IoU threshold of 0.5 and set the classification confidence threshold at 0.3.
@@
Table 1: Effect of TTC system on various 3D detectors. TTC system effectively improves the test-time performance of offline-trained detectors during online inference without any extra training.
Method Backbone Type mAP (%) $\uparrow$ EDS (%) $\uparrow$
MonoDETR [67] R101 Monocular 37.9 38.2
TTC-MonoDETR 42.6 (+5.7) 43.2 (+5.0)
MV2D [57] R50 Multiview 38.9 38.3
TTC-MV2D 51.0 (+12.1) 51.0 (+12.7)
Sparse4Dv2 [31] R50 Multiview + Temporal 38.8 37.6
TTC-Sparse4Dv2 53.5 (+14.7) 52.1 (+14.5)
BEVFormer [27] R101 BEV + Temporal 36.5 35.8
TTC-BEVFormer 47.8 (+11.3) 47.2 (+11.4)
BEVFormerV2-t8 [62] R50 BEV +Temporal 39.6 38.9
TTC-BEVFormerV2-t8 51.6 (+12.0) 51.0 (+12.1)
RayDN [32] R50 BEV + Temporal 39.6 38.7
TTC-RayDN 51.7 (+12.1) 50.8 (+12.1)
StreamPETR [54] V2-99 BEV + Temporal 39.7 39.1
TTC-StreamPETR 52.3 (+12.6) 51.5 (+12.4)
@@
Point 41.6 40.8 17.8 19.4
Box 44.3 43.7 19.2 21.7
Visual 42.6 42.0 18.3 21.6
MonoDETR 31.4 31.6 0.0 0.0
TTC-MonoDETR 40.2 40.1 11.0 14.4
$\Delta$ +8.8 +8.5 +11.0 +14.4
@@
Point 38.0 36.9 33.5 36.6
Box 41.2 40.0 36.7 39.6
Visual 39.2 38.4 34.6 38.5
MonoDETR 17.6 16.1 2.4 7.3
TTC-MonoDETR 29.0 27.2 23.2 28.9
$\Delta$ +11.4 +11.1 +20.8 +21.6
@@
Point 35.2 35.6 11.9 15.0
Box 38.5 38.8 14.7 16.9
Visual 35.4 36.0 11.2 15.6
MonoDETR 28.3 29.6 0.0 0.0
TTC-MonoDETR 34.8 36.1 8.5 13.6
$\Delta$ +6.5 +6.5 +8.5 +13.6
@@
Point 39.0 38.2 30.5 30.7
Box 42.6 41.5 33.4 33.6
Visual 39.8 39.4 29.4 30.8
MonoDETR 34.5 34.7 25.2 26.6
TTC-MonoDETR 39.9 40.0 29.7 31.3
$\Delta$ +5.4 +5.3 +4.5 +4.7
@@ 4.2 Main Result
Effectiveness of TTC in Test-time Error Correction. Test-time error correction ability without re-training is the core capability of TTC system. We verify this by incorporating traditional offline-trained 3D detectors into TTC system and compare the performance without test-time parameter updates. For thorough verification, we select various offline-trained 3D detectors, including monocular [67], multi-view [57, 31], and BEV ones [27, 62, 54, 32]. As shown in Table 1, the TTC system substantially improves the test-time performance of offline-trained 3D detectors, e.g., 11.4% and 12.4% EDS improvements on BEVFormer and StreamPETR, without requiring additional training during inference. These demonstrate the effectiveness of the TTC system in instantly correcting test-time errors during online inference.
@@
Effectiveness of TTC in Out-of-Training Scenarios. Table 2 presents results to validate the TTC system in challenging and extreme scenarios. In these experiments, we base TTC system on MonoDETR for its simple monocular setting. “Point”, “Box”, and “Visual” in Table 2 are TTC-MonoDETR using point, box, or image patch of target objects in the input image as input prompts. These are performance upper bounds as they receive human feedback in every frame.
@@
Table 2(a) evaluates the effectiveness of TTC system in rectifying detection errors on distant objects (beyond 30m). During training, 80% of the far-away annotations are removed, resulting in a 3D detector with poor long-range performance (0.0% mAP and EDS). Powered by the TTC system, the performance on these distant objects is instantly improved to 11.0% mAP and 14.4% EDS, without any extra training. The experiment is then extended to all vehicles in the nuScenes dataset, as shown in Table 2(b). The TTC achieves impressive performance gains of 20.8% mAP and 21.6% EDS.
@@
Furthermore, in Table 2(c) and Table 2(d), we evaluate the TTC system in two challenging scenarios: encountering unseen objects not present in the training data; handling domain shifts, such as transitioning from sunny to rainy or nighttime conditions. In Table 2(c), the TTC system shows effectiveness in successfully detecting novel objects not labeled in the training set, achieving 8.5% mAP and 13.6% EDS on these novel objects, significantly improving the offline-trained baseline with 0.0% mAP and EDS. In Table 2(d), we find TTC also works well for online error rectification when driving into scenarios with domain shifts, providing 4.5% mAP and 4.7% EDS improvements.
@@
Regarding qualitative results, Figure 5 further illustrates a case of the zero-shot capability in the real-world. Despite being trained solely on the nuScenes dataset, TTC-MonoDETR can detect a Deer using a visual prompt of another deer from a different viewpoint, which is extremely challenging for traditional offline detectors. More qualitative examples can be found in supplementary material Appendix E.
@@ 4.3 Robustness of Visual Prompts
As a crucial component of the continuous test-time error correction system, we conduct a series of ablation studies on visual prompts in this section. We base TTC system on MonoDETR and validate its robustness concerning prompts obtained from the Internet or those from different moments.
@@
Car MonoDETR 3.2 9.1
TTC-MonoDETR 20.9 (+17.7) 27.5 (+18.4)
Bus MonoDETR 0.4 4.4
TTC-MonoDETR 17.4 (+17.0) 20.6 (+16.2)
@@
Robustness over Web-derived Visual Prompts. Visual prompts can be arbitrary imagery views of target objects and can be from any image source. For example, we can use images sourced from the Internet as visual prompts. We fix the prompt buffer with visual prompts from the Internet during inference, and assess the effectiveness of handling visual prompts with diverse styles. We employ the model from Table 2(b) (TTC-MonoDETR trained with 20% labels of vehicles), and select 12 car and 6 bus images from the Internet, which resemble those in nuScenes dataset, as visual prompts for online correction. We show them in Figure 6. As listed in Table 3, TTC system demonstrates strong online correction capabilities though with limited annotations. Even with prompts sourced from the Internet with various styles and poses, TTC system still improves the offline-trained baseline by 18.4% EDS on “Car” objects under this extremely challenging setting. This further underscores the robustness of TTC system and highlights its potential to address the long-tail challenges in real-world scenarios.

### v3

@@ 3.3 Visual Prompt Buffer
Visual prompt buffer is a queue that stores visual prompts of missed objects during online inference. In the default setting, these visual prompts are the 2D regions of risky objects in previous frames. The TTC system is also flexible enough to allow freely uploading visual prompts, which can be either image content from the current scene or customized images sourced from the Internet based on road descriptions or user-specified requests. This versatility makes the TTC system applicable to a wide spectrum of scenarios.
Dequeue. To prevent unbounded growth, we adopt a dequeue policy that combines score-based eviction and redundancy pruning. During training, we mix in negative visual prompts—patches that do not correspond to any target in the current frame (e.g., background crops or off-timestamp instances)—and supervise the model to assign them low confidence. This calibration causes non-existent or contradictory prompts to be naturally down-weighted at test time, enabling confidence-based removal. Concretely, we first filter out prompts whose associated detections consistently fall below a confidence threshold. We then suppress near-duplicates by measuring intersection-over-union (IoU) between predictions and removing redundant prompts. This dynamic maintenance balances latency and accuracy by adaptively incorporating informative prompts while pruning stale or duplicate ones online.
@@ 4 Experiments
We proceed to evaluate our TTC system from two key aspects:
@@ 4.1 Experimental Setup
Dataset. Experiments are done on the nuScenes dataset [nuscenes2019] with 1,000 autonomous driving sequences, one of the most popular datasets for autonomous driving research.
@@
Tasks. Experiments are conducted in two aspects to answer the questions above. The first experiment aims to test the TTC system in boosting the performance of offline-trained 3D detectors via online error rectification. We conduct this verification by applying TTC system to various established offline 3D detectors [zhang2022monodetr, li2022bevformer, yang2023bevformerV2, liu2024ray, wang2023streampetr, sparse4dv2, wang2023object].
@@
Then, we validate the TTC system to correct detection errors in out-of-training-distribution scenarios. To conduct a thorough quantitative analysis of this aspect, we established four tasks under different settings: (a.) discarding 80% labels of distant objects farther than 30m during training, and test the error corrections for detecting distant objects; (b.) discarding 80% labels of instances labeled as vehicle, including “car”, “truck”, “C.V.”, “bus”, and “trailer”, and test the error correction in vehicles; (c.) discarding all annotations of class “truck” and “bus”, and test the zero-shot ability of TTC on those discarded classes; (d.) discarding all training data of the scenario of “Nighttime” and “Rainy”, and test the improvements of TTC on scenarios with domain gap. For these tasks, we base TTC system on the monocular algorithm, MonoDETR [zhang2022monodetr], and test it under each setting separately, as the monocular setting demonstrates the most general applicability. Based on TTC-MonoDETR, we also present qualitative results to show the potential of TTC to address corner cases in challenging real-world scenarios through auxiliary feedback during test time. Through these experiments, we demonstrate TTC as an effective system to adapt offline 3D detectors to challenging scenarios.
@@
Entity Detection Score. As we focus on enabling existing 3D detectors to rectify missing detections, both in-distribution and out-of-distribution, we adopt a class-agnostic setting. We remove all class annotations of 3D objects during training, treating all objects as entities [kirillov2023segany, qi2022openworld]. For evaluation, we use the Entity Detection Score (EDS), a class-agnostic version of the nuScenes Detection Score (NDS).
[…]
@@
TABLE 1: Effect of TTC on various 3D detectors. TTC improves test-time performance of offline-trained detectors during online inference without extra training. To compare directly with the baselines’ official scores, we set the confidence threshold to 0.01. For false-alarm analysis, we report FP/frame and FN/frame.
Method Backbone Type mAP (%) $\uparrow$ EDS (%) $\uparrow$ FP/frame@0.01 $\downarrow$ FN/frame@0.01 $\downarrow$
MonoDETR [zhang2022monodetr] R101 Monocular 42.6 50.1 49.9 1.2
TTC-MonoDETR 45.0 (+2.4) 51.7 (+1.6) 49.9 0.9
MV2D [wang2023object] R50 Multiview 43.0 52.8 32.2 2.0
TTC-MV2D 52.7 (+9.7) 57.2 (+4.4) 32.2 0.8
Sparse4Dv2 [sparse4dv2] R50 Multiview + Temporal 44.4 53.2 42.7 1.9
TTC-Sparse4Dv2 54.2 (+9.8) 57.8 (+4.6) 42.7 0.7
BEVFormer [li2022bevformer] R101 BEV + Temporal 41.1 50.2 34.2 1.9
TTC-BEVFormer 50.2 (+9.1) 54.4 (+4.2) 34.4 0.8
BEVFormerV2-t8 [yang2023bevformerV2] R50 BEV + Temporal 45.6 54.1 42.2 1.7
TTC-BEVFormerV2-t8 54.9 (+9.3) 58.6 (+4.5) 42.3 0.6
RayDN [liu2024ray] R50 BEV + Temporal 44.9 53.6 39.8 1.8
TTC-RayDN 54.4 (+9.5) 58.1 (+4.5) 39.8 0.7
StreamPETR [wang2023streampetr] V2-99 BEV + Temporal 45.0 53.8 42.3 1.8
TTC-StreamPETR 51.6 (+6.6) 55.6 (+1.8) 42.4 0.7
@@
Point 41.6 52.5 17.8 35.0
Box 44.3 54.3 19.2 36.3
Visual 42.6 53.0 18.3 36.1
MonoDETR 31.4 47.6 0.0 0.0
TTC-MonoDETR 39.6 52.1 10.1 27.9
$\Delta$ +8.2 +4.5 +10.1 +27.9
@@
Point 38.0 48.2 33.5 52.8
Box 41.2 50.3 36.7 54.4
Visual 39.2 47.7 34.6 51.8
MonoDETR 17.6 29.0 2.4 36.0
TTC-MonoDETR 28.5 35.4 22.4 44.2
$\Delta$ +10.9 +6.4 +20.0 +8.2
@@
Point 35.2 48.7 11.9 33.3
Box 38.5 50.9 17.0 36.6
Visual 35.4 47.3 11.2 33.3
MonoDETR 28.3 47.0 0.0 0.0
TTC-MonoDETR 34.6 49.0 7.2 30.8
$\Delta$ +6.3 +2.0 +7.2 +30.8
@@
Point 39.0 49.4 30.5 44.3
Box 42.6 51.8 33.4 46.3
Visual 39.8 48.6 29.4 43.0
MonoDETR 34.5 48.4 25.2 43.1
TTC-MonoDETR 39.5 51.8 29.0 44.5
$\Delta$ +5.0 +3.4 +3.8 +1.4
@@ 4.2 Main Result
Effectiveness of TTC in Test-time Error Correction. Test-time error correction ability without re-training is the core capability of the TTC system. We verify this by incorporating traditional offline-trained 3D detectors into the TTC system and compare the performance without test-time parameter updates. For thorough verification, we select various offline-trained 3D detectors, including monocular [zhang2022monodetr], multi-view [wang2023object, sparse4dv2], and BEV ones [li2022bevformer, yang2023bevformerV2, wang2023streampetr, liu2024ray]. The online feedback is derived from the human clicks which is then converted into visual prompts. The feedback interval is set to 5 frames. As shown in Table˜1, the TTC system substantially improves the test-time performance of offline-trained 3D detectors, e.g., 9.1% and 6.6% mAP improvements on BEVFormer and StreamPETR, without requiring additional training after deployment. These demonstrate the effectiveness of the TTC system in instantly correcting test-time errors during online inference.
@@
Effectiveness of TTC in Out-of-Training Scenarios. Table˜2 presents results to validate the TTC system in challenging and extreme scenarios. In these experiments, we base TTC system on MonoDETR for its simple monocular setting. “Point”, “Box”, and “Visual” in Table˜2 are TTC-MonoDETR directly using point, box, or image patch of target objects in the input image as input prompts, which are impractical unless automated annotation scenario. These are performance upper bounds as they receive human feedback in every frame. Compared with the baseline MonoDETR, “TTC-MonoDETR” in the table obtains visual prompts of missing objects from user at a feedback interval of 5 frames, enabling continuous correction without any test-time weight updates.
@@
Table˜II(a) evaluates the effectiveness of TTC system in rectifying detection errors on distant objects (beyond 30m). During training, 80% of the far-away annotations are removed, resulting in a 3D detector with poor long-range performance (0.0% mAP and EDS). Powered by the TTC system, the performance on these distant objects is instantly improved to 10.1% mAP and 27.9% EDS, without any extra training. The experiment is then extended to all vehicles in the nuScenes dataset, as shown in Table˜II(b). The TTC achieves impressive performance gains of 20.0% mAP and 8.2% EDS.
@@
Furthermore, in Table˜II(c) and Table˜II(d), we evaluate the TTC system in two challenging scenarios: encountering unseen objects not present in the training data; handling domain shifts, such as transitioning from sunny to rainy or nighttime conditions. In Table˜II(c), the TTC system shows effectiveness in successfully detecting novel objects not labeled in the training set, achieving 7.2% mAP and 30.8% EDS on these novel objects, significantly improving the offline-trained baseline with 0.0% mAP and EDS. In Table˜II(d), we find TTC also works well for online error rectification when driving into scenarios with domain shifts, providing 3.8% mAP and 1.4% EDS improvements.
@@
Regarding qualitative results, Figure˜5 further illustrates a case of the zero-shot capability in the real world. Despite being trained solely on the nuScenes dataset, TTC-MonoDETR can detect a Deer using a visual prompt of another deer from a different viewpoint (simulating an internet image search for road description “deers might appear on this road”), which is challenging for traditional offline detectors. More qualitative examples can be found in Appendix Appendix˜C.
@@ 4.3 Robustness of Visual Prompts
As a crucial component of the continuous test-time error correction system, we conduct a series of experiments on visual prompts in this section. Specifically, we analyze the robustness of visual prompts against: (i) variations in prompt sources and styles, (ii) changes in viewpoints across frames, (iii) spatial perturbations in provided prompts, (iv) reduced feedback frequency, (v) noisy (off-topic) visual prompts, and (vi) rapidly changing prompts.
@@
Car MonoDETR 3.2 37.5
TTC-MonoDETR 20.9 (+17.7) 45.3 (+7.8)
Bus MonoDETR 0.4 28.1
TTC-MonoDETR 17.4 (+17.0) 40.2 (+12.1)
@@
Robustness over Web-derived Visual Prompts. Visual prompts can be arbitrary imagery views of target objects and can be from any image source and styles. For example, we can use images sourced from the Internet as visual prompts. We fix the prompt buffer with visual prompts from the Internet during inference, and assess the effectiveness of handling visual prompts with diverse styles. We employ the model from Table˜II(b) (TTC-MonoDETR trained with 20% labels of vehicles), and select 12 car and 6 bus images from the Internet, which resemble those in the nuScenes dataset, as visual prompts for online correction. We show them in Figure˜6. As listed in Table˜3, TTC system demonstrates strong online correction capabilities. Even with prompts sourced from the Internet with various styles and poses, TTC system still improves the offline-trained baseline by 17.7% mAP on “Car” objects under this extremely challenging setting. This further underscores the robustness of TTC system and highlights its potential to address the long-tail challenges in real-world scenarios.